# The Software Diversity Card: A Framework for Reporting Diversity in Software Projects

### v1

@@ Abstract
The interest and concerns about diversity in software development have soared in recent years. Reporting diversity-related aspects of software projects can increase user trust and help regulators evaluate potential adoption. Furthermore, recent directives around AI are beginning to require diversity information in the development of AI products, indicating the growing interest of public regulators in it. Despite this importance, current documentation assets in software development processes frequently overlook diversity in favor of technical features, partly due to a lack of tools for describing and annotating diversity.
This work introduces the Software Diversity Card, a comprehensive framework for reporting diversity-related aspects of software projects. The card is designed to profile the different types of teams involved in developing and governing software projects (including the final user groups involved in testing), and the software adaptations for specific social groups. To encourage its adoption, we provide a diversity modeling language, a toolkit for generating the cards using such language, and a collection of real-world examples from active software projects. Our proposal can enhance diversity practices in software development (e.g., through open-source projects like the CONTRIBUTING.md file), support public administrations in software assessment, and help businesses promote diversity as a key asset.
@@ 1 Introduction
Learning about the people engaged in software product creation can help users better understand the software they use, thereby increasing trust. For example, a period tracker app designed for women may inspire greater user confidence if the developers and testers include women sharing similar experiences and needs. Similarly, when developing software for specific social groups, it is essential to consider their context and unique characteristics. For example, Bangladeshi fisherfolk [10] faced significant barriers in adopting new software technologies due to factors such as low digital literacy, varying language styles, and their community’s traditional knowledge transmission methods. These kinds of barriers can be mitigated by gaining a better understanding of the social group’s context and specific needs.
Concerns about sharing diversity information are also increasing in other domains, impacting public regulatory bodies worldwide. For instance, influential documentation frameworks in the Machine Learning (ML) field stress the need to document the diversity of actors involved in the data creation processes [11, 12, 13]. These recommendations have been incorporated into recent regulations on artificial intelligence, such as the European AI Act [^1] AI Act required documentation - Annex IV: https://www.euaiact.com/annex/4 and the US Bill of Rights [^2] BoR homepage: https://bidenwhitehouse.archives.gov/ostp/ai-bill-of-rights, and are expected to become a fundamental practice in developing future responsible AI applications.
@@
Despite the undeniable benefits of reporting diversity-related aspects of software projects, there is neither current standard documentation practices nor concrete reporting technique for this purpose. Current documentation assets in software engineering focus on the technical features of software, often overlooking the social and diversity aspects.
@@
In the remainder of this document, Section 2 presents the background and related work on diversity and human aspects in software development, while Section 3 empirically evaluates the current reporting diversity practices in open-source communities motivating the need for our proposal. Then, Section 4 presents the main concepts and structure of the card. Section 5 describes the development of the proposal, introducing a diveristy domain-specific language and a toolkit (including a language plugin [^3] Language plugin repository: https://github.com/SOM-Research/SoftwareDiversityCard and a form-based web editor [^4] Online demo of the web editor: https://huggingface.co/spaces/JoanGiner/SoftwareDiversityCard-Generator) that supports developers in creating diversity cards for their projects. In Section 6, we present two running examples of real-world software projects described using our proposal. Finally, in Section 7, we discuss the challenges and open research lines on reporting diversity information in software projects.
@@ 2 State of the art
This section reviews previous research on diversity and inclusion in software development, identifying the dimensions and teams where diversity plays an important role in software projects. This will serve as the foundation for our work. To that end, we begin by reviewing works that focus on the development team, then on other types of participants in software projects (i.e., non-coding roles), and finally on the importance of diversity for the end users and the project adaptations for specific populations. Finally, we conduct a brief review of ML documentation frameworks, assessing the value of profiling those involved in project governance.
2.1 Software developers diversity and inclusion
The intersection of politics and economics is fundamental in discussions of equity and inclusion in computer science, as technological advancements and software projects are influenced by social, economic, and political factors. Relevant works [1, 2] in Software Developers’ Diversity and Inclusion (SDDI) have pointed out that the economic structures influence those who have access to technology, whose perspectives are prioritized in software development and whose biases become embedded in digital systems. As a relevant aspect of our proposal, these works have defined a set of essential dimensions such as gender, ethnicity/race, and disabilities, among others, that will serve as the foundation for our profiles. Moreover, another relevant research [21] has conducted an empirical study on diversity awareness in the software engineering research community, identifying similar dimensions.
@@ 2.1 Software developers diversity and inclusion
In terms of gender diversity, a recent study [22] shows that lack of gender diversity remains an issue, since they observed less than 10% of female-gender developers in the open-source projects analyzed. In response to this situation, Oliveira et al. [7] have studied the path of women in software engineering, detecting persistent challenges in their career development. In the same line, Gila et al. [23] described the effects of gender diversity on team performance based on personality types that show that female-gender developers looked uneasy in male-dominant teams. In contrast to this, and motivating the field of research, recent findings [24] show the clear benefits of gender-diversity in developing software products in other scientific fields [3].
@@
Also, regarding the race/ethnicity of the development team, Nadri et al. [25] presents a novel qualitative analysis of non-merged pull requests and classifies the reasons why the pull requests were not merged based on four perceived racial identities: Asian/Pacific Islander, Black, Hispanic, and White. The authors found that perceptibly Black and Asian/Pacific developers had pull requests rejected at higher rates than those of perceptibly White developers. Other works as [26] examine how geographic location associates with pull request submission and acceptance rates in open source projects. The authors found that most closed pull requests came from developers in the United States, the United Kingdom, and Germany, countries with a high HDI [^5] The Human Development Index (HDI) is a measure of health, education, income and living conditions in a each country, compiled by the United Nations Development Programme..
@@
Disabilities have also been a focus of interest. For instance, recent research in the field has focused on creating more inclusive and diverse teams and integrating socially protected groups such as neurodivergent people. For example, Silva et al. [27] identifies the main techniques in Agile Methodologies to promote the inclusion of people with disabilities in development teams, and Liebel et al. [28] has proposed a set of recommendations to the industry to include developers suffering from ADHD (Attention-Deficit/Hyperactivity Disorder). Furthermore, answering the need for intersectional analysis of diversity, Prana et al. [29] have analyzed the intersection of gender and region origin in the research community. This study considered the consistency of gender inequalities across different regions.
@@ 2.2 Diversity comprises a wider definition of participants
However, a wider range of profiles often participate in software projects. Recent research has highlighted the relevant role of these profiles in software communities [15]: individuals commenting and reacting in open-source communities, pull request reviewers, bug reporters, translators, or developer advocates. In addition, recent studies about open-source communities [30] have highlighted the toxicity in these communities, motivating the need to profile communities if we want to avoid toxic behaviors.
In this sense, Paul et al. [31] looked at code reviews, detecting that male reviewers were more likely to provide negative comments to female contributors. In contrast, Zolduoarrati et al. [32] analyzed Stack Overflow archival data over 11 years, showing that female contributions tend to be more cooperative, supportive and collective and that women are more willing to share knowledge than their male counterparts. Another profile that has raised concerns in terms of diversity is that of modelers, and recent works [33] in the model-driven community also focused on the human aspects of modelers, emphasizing the modelers’ diversity and highlighting the relevance of their background in modeling decisions.
@@ 2.3 Diversity and human aspects impacting software end-users
Diversity and the human aspects of end-users play a critical role in software projects, as they directly impact usability, accessibility, and overall user satisfaction. Research has increasingly highlighted the importance of inclusiveness in software development, emphasizing that neglecting diverse user characteristics, limitations, and abilities can lead to dissatisfaction and exclusion. Studies such as those by [34] and [35] propose taxonomies and analytical frameworks to better understand human-centric issues in software engineering. A key challenge is that inclusiveness is often overlooked during the development process, resulting in software that fails to accommodate the full range of human diversity [36, 37]. For instance, by integrating user feedback throughout the life-cycle, developers can ensure that applications serve a broader audience while minimizing the risk of bias and accessibility barriers [38].
Efforts to adapt software to specific populations have already been made in the research community. The concept of Adaptive User Interfaces (AUIs) has gained traction in research [39, 40], offering solutions that personalize interfaces for different user groups, including individuals with chronic conditions [41]. Similarly, crowd-testing has emerged as a valuable approach in software development, enabling diverse users to participate in the testing phase and provide real-world insights into software usability. Studies on crowd testers, such as those by [42, 9], emphasize the importance of selecting an appropriate tester community to ensure a representative evaluation of software.
@@ 2.4 Diversity in machine learning documentation frameworks
Recent documentation and reporting frameworks in the Machine Learning (ML) field, such as Datasheets for Datasets [11], Data Statements [43], and Data Cards [12], have become crucial to ensuring transparency, accountability, and fairness of these systems. These frameworks aim to capture the various aspects of data collection, model development, and evaluation while explicitly addressing biases, ethical considerations, and the societal impact of ML technologies. By incorporating diversity-related metadata into documentation, stakeholders can better assess whether ML systems fairly represent different demographic groups, mitigate potential biases, and adhere to inclusive algorithmic decision-making principles. Public regulatory initiatives such as the European AI Act [^6] AI Act required documentation - Annex IV: https://www.euaiact.com/annex/4 and the AI Bill of Rights [^7] Bill’s official website: https://www.whitehouse.gov/ostp/ai-bill-of-rights have started to adopt them, showing the increasing interest of public administrations in diversity.
One of the relevant points of these frameworks is the wider definition of participants behind an ML system, beyond the development team. Such frameworks aim to profile those individuals involved in gathering and annotating the data used to train the ML algorithms, as well as those in charge of governing the project, showing the relevance of these groups in decisions that may affect the final behavior of the ML models. In relation to our work, ML systems are, in fact, software systems. There is a path to study the integration of this ML-specific diversity needs (such as profiling data labellers) with broader software documentation proposals as we propose in this work. On the other hand, these works show the relevance of assessing also those who govern the projects, a case specially relevant for open-source communities, with examples of relevant open-source communities as Drupal [^8] Drupal’s published board association profiles. or Vue.js [^9] Vue.js published governance wiki: https://github.com/vuejs/governance who disclose publicly their governing participants and policies.
@@ 3.1 Study method
The search for diversity documentation evidence has focused on five main groups extracted from the review of the previous section, and further presented in Section 4. Specifically: (A1) the development team, and whether they mention profile aspects about the developers; (A2) non-coding contributors, and whether they explain non-coding roles; (A3) tests with potential users, and whether they mention labor force or report platforms; (A4) usage context, including the target population and possible specific adaptations; and (A5) governance, including information on funders. Areas A1 and A2 are motivated by the works on diversity for developers and non-coding roles (cf. Sections 2.1 and 2.2), while A3 and A4 are inspired by the need for specifying human aspects (cf. Section 2.3). The works on ML documentation (cf. Section 2.4) prompted A5. For each area, we reported whether the OSS project included (or not) such information.
@@ 1. Does the text mention the development team explicitly?
To perform the analysis and search for diversity documentation, we built a tool based on Large Language Models (LLMs) [^12] Tool’s repository: https://github.com/SOM-Research/Analyzer-Diversity-Card. In particular, we relied on GPT-4o-mini. The use of LLMs for analyzing text has demonstrated their effectiveness, for instance, for sentiment analysis [45, 46], text classification [47, 48], studying social interactions [49], and detecting ethical deviations [50]. In our context, the use of LLMs eliminates the need for extensive preprocessing, enabling a more streamlined and direct analysis of the text. Our tool receives the extracted files as input and performs a prompt-driven search of diversity information for each area. Each prompt instructs the LLM to look for evidence in a specific area and includes some examples (i.e., few-shot prompting). Listing 1 shows an excerpt of one of the five prompts used, in particular, the one addressing A1131313The complete list of prompts used in the study can be found at https://doi.org/10.5281/zenodo.14981689. To validate the tool, we manually checked the evidence found for each area for a sample of 50 random projects from our dataset, obtaining a 92.54% agreement between our assessment and the tool output.

### v2

@@ Abstract
Context: Interest in diversity in software development has significantly increased in recent years. Reporting on diversity in software projects can enhance user trust and assist regulators in evaluating adoption. Recent AI directives include clauses that mandate diversity information during development, highlighting the growing interest of public regulators. However, current documentation often neglects diversity in favor of technical features, partly due to a lack of tools for its description and annotation.
Objectives: This work introduces the Software Diversity Card, a structured approach to document and share diversity-related aspects within software projects. It aims to profile the various teams involved in software development and governance, including user groups in testing and software adaptations for diverse social groups.
Methods: We performed a literature review on diversity and inclusion in software development and an analysis of 1,000 top-starred Open Source Software (OSS) repositories in GitHub to identify diversity-related information. Moreover, we present a diversity modeling language, a toolkit for generating the cards using that language, and a study of its application in two real-world software projects.
Results: Despite the growing awareness of diversity in the research community, our analysis found a notable lack of diversity reporting in OSS projects. Applying the card to real-world examples highlighted challenges like balancing anonymity with transparency, managing sensitive data, and ensuring authenticity.
Conclusion: We believe that our proposal can enhance diversity practices in software development, support public administrations in software assessment, and help businesses promote diversity as a key asset.
@@ 1 Introduction
Learning about the people engaged in software product creation can help users better understand the software they use, thereby increasing trust [10, 11]. For example, a period tracker app designed for women may inspire greater user confidence if the developers and testers include women sharing similar experiences and needs. Similarly, when developing software for specific social groups, it is essential to consider their context and unique characteristics. For example, Bangladeshi fisherfolk faced significant barriers in adopting new software technologies due to factors such as low digital literacy, varying language styles, and their community’s traditional knowledge transmission methods. These kinds of barriers can be mitigated by gaining a better understanding of the social group’s context and specific needs [12].
Concerns about sharing diversity information are also increasing in other domains, impacting public regulatory bodies worldwide. For instance, influential documentation frameworks in the Machine Learning (ML) field stress the need to document the diversity of actors involved in the data creation processes [13, 14, 15]. These recommendations have been incorporated into recent regulations on artificial intelligence, such as the European AI Act [^1] EU AI Act required documentation - Annex IV: https://www.euaiact.com/annex/4 and the US AI Bill of Rights [^2] US AI Bill of Rights: https://bidenwhitehouse.archives.gov/ostp/ai-bill-of-rights, and are expected to become a fundamental practice in developing future responsible AI applications.
@@
Despite the clear benefits of reporting diversity-related aspects of software projects, there are currently no standardized documentation practices or concrete reporting techniques for this purpose. Existing documentation in software engineering primarily emphasizes technical features, often overlooking social and diversity considerations, while current machine learning documentation frameworks tend to focus on data assets, frequently neglecting the broader complexities of software projects.
@@
In the remainder of this document, Section 2 presents the background and related work on diversity and human aspects in software development, while Section 3 empirically evaluates the current reporting diversity practices in open-source communities motivating the need for our proposal. Then, Section 4 presents the main concepts and structure of the card. Section 5 describes the development of the proposal, introducing a diversity domain-specific language and a toolkit (including a language plugin [^3] Language plugin repository: https://github.com/SOM-Research/SoftwareDiversityCard and a form-based web editor [^4] Online demo of the web editor: https://huggingface.co/spaces/JoanGiner/SoftwareDiversityCard-Generator) that supports developers in creating diversity cards for their projects. In Section 6, we present two running examples of real-world software projects described using our proposal. Finally, in Section 7, we discuss the challenges and limits of our study, while in Section 8, we present the conclusion and open research lines on reporting diversity in software projects.
@@ 2 State of the art
This section reviews prior research on diversity and inclusion in software development. We begin by examining the dimensions that previous studies have used to characterize diversity, which provides the foundation for our own characterization of teams in software projects. Next, we explore research focused on diversity within different types of teams in software projects, including both development teams and non-coding contributors. We then discuss the significance of diversity for end users and how projects adapt to meet the needs of specific populations. Finally, we provide a brief overview of emerging machine learning documentation frameworks, highlighting their opportunities and limitations in software projects.
2.1 Characterizing diversity in software
Relevant works on Software Developers’ Diversity and Inclusion (SDDI) have shown that economic and political structures shape who gains access to technology, whose perspectives are prioritized in software development, and whose biases become embedded in digital systems [1, 2]. These findings highlight the importance of examining the profiles of those involved in the complex ecosystem of software projects. A central contribution of prior work has been the identification of a set of dimensions for studying and characterizing diversity within software engineering.
In this regard, Dutta et al. [23] analyzed 79 research papers and 105 participant studies in the SDDI community, identifying the most common dimensions used to describe individuals involved in software projects. These include gender, age, ethnicity/race, socioeconomic status, and sexual orientation, as well as professional attributes such as tenure, job title, language expertise, and education level. The selection of these dimensions is also reflected in recent influential studies [1] and community-driven initiatives, such as special issues on the diversity crisis in software development [2], that underscored the need to to account also for disabilities and intersectionality during the study of diversity.
Together, these efforts establish a baseline for characterizing diversity in software projects. While this characterization is dynamic and may evolve as new societal needs emerge, it provides the foundation for our selection of dimensions in this study, shown in Figure 2. To underscore the relevance of these dimensions, the following section reviews prior work that has examined diversity across them, with a focus on both development teams first and then on other types of teams involved directly and indirectly in software projects.
2.2 Diversity on software development teams
The study of the composition of open-source projects has been a focus of interest of the research community. One of the seminal studies is the proposal of the onion model [24], which studies the software project’s social structure as concentric layers of contributors. The model highlights core contributors at the center, surrounded by peripheral and occasional participants, illustrating roles, responsibilities, and patterns of collaboration within the community. An excellent case study illustrating the structure of real-world open-source software projects is the GNOME project [25] and subsequent studies on the composition and inherent nature of the contributors [26, 27]. Going beyond the composition, posterior studies [28] explored how both intrinsic and extrinsic factors shape contributor engagement, relating to the focus of interest of this work.
@@ 2.2 Diversity on software development teams
The composition of the development team has also been studied regarding the gender dimensions. Studies show that the lack of gender diversity remains an issue [29], since they observed less than 10% of female developers in the open-source projects analyzed. In response to this situation, Oliveira et al. [7] have studied the path of women in software engineering, detecting persistent challenges in their career development. In the same line, Gila et al. [30] described the effects of gender diversity on team performance based on personality types that show that female-gender developers looked uneasy in male-dominant teams. In contrast to this, and motivating the field of research, recent findings [31] show the clear benefits of gender-diversity in developing software products in other scientific fields [3].
@@
Also, regarding the race/ethnicity of the development team, Nadri et al. [32] presents a novel qualitative analysis of non-merged pull requests and classifies the reasons why the pull requests were not merged based on four perceived racial identities: Asian/Pacific Islander, Black, Hispanic, and White. The authors found that perceptibly Black and Asian/Pacific developers had pull requests rejected at higher rates than those of perceptibly White developers. Other works as [33] examine how geographic location associates with pull request submission and acceptance rates in open source projects. The authors found that most closed pull requests came from developers in the United States, the United Kingdom, and Germany, countries with a high HDI [^5] The Human Development Index (HDI) is a measure of health, education, income and living conditions in a each country, compiled by the United Nations Development Programme.. Finally, several works, such as [34, 35], mixed gender and nationality to investigate their relation within GitHub teams, analyzing how these factors vary across software projects and how demographic diversity can influence participation patterns and collaboration dynamics within software development teams.
@@
Disabilities have recently arisen as a focus of interest. For instance, recent research in the field has focused on creating more inclusive and diverse teams and integrating socially protected groups such as neurodivergent people. For example, Silva et al. [36] identify the main techniques in Agile Methodologies to promote the inclusion of people with disabilities in development teams, and Liebel et al. [37] have proposed a set of recommendations to the industry to include developers suffering from ADHD (Attention-Deficit/Hyperactivity Disorder). Furthermore, answering the need for intersectional analysis of diversity, Prana et al. [38] have analyzed the intersection of gender and region origin in the research community. This study considered the consistency of gender inequalities across different regions.
@@ 2.3 Diversity comprises a wider definition of participants
However, a wider range of profiles often participate in software projects. Recent research has highlighted the relevant role of these profiles in software communities [17]: individuals commenting and reacting in open-source communities, pull request reviewers, bug reporters, documenters, translators, or developer advocates. In addition, recent studies about open-source communities [39] have highlighted the toxicity in these communities, motivating the need to profile communities if we want to avoid toxic behaviors.
In this sense, Paul et al. [40] looked at code reviews, detecting that male reviewers were more likely to provide negative comments to female contributors. In contrast, Zolduoarrati and Licorish [41] analyzed Stack Overflow archival data over 11 years, showing that female contributions tend to be more cooperative, supportive and collective and that women are more willing to share knowledge than their male counterparts. Roles regarding the creation and maintenance of the documentation for software projects [42] and the translation of this documentation [43] have also been studied pointing out their relevance inside a software project. Another profile that has raised concerns in terms of diversity is that of modelers, and recent works [44] in the model-driven community also focused on the human aspects of modelers, emphasizing the modelers’ diversity and highlighting the relevance of their background in modeling decisions.
@@ 2.4 Diversity and human aspects impacting software end-users
Diversity and the human aspects of end-users play a critical role in software projects, as they directly impact usability, accessibility, and overall user satisfaction. Seminal works on socio-technical systems have highlighted the relevant roles of the social context around software systems [45], and have increasingly highlighted the importance of inclusiveness in software development, emphasizing that neglecting diverse user characteristics, limitations, and abilities can lead to dissatisfaction and exclusion. Studies such as those by [46] and [47] propose taxonomies and analytical frameworks to better understand human-centric issues in software engineering. A key challenge is that inclusiveness is often overlooked during the development process, resulting in software that fails to accommodate the full range of human diversity [48, 49]. For instance, by integrating user feedback throughout the life-cycle, developers can ensure that applications serve a broader audience while minimizing the risk of bias and accessibility barriers [50].
Efforts to adapt software to specific populations have already been made in the research community. The concept of Adaptive User Interfaces (AUIs) has gained traction in research [51, 52], offering solutions that personalize interfaces for different user groups, including individuals with chronic conditions [52]. Similarly, crowd-testing has emerged as a valuable approach in software development, enabling diverse users to participate in the testing phase and provide real-world insights into software usability. Studies on crowd testers, such as [9, 53], emphasize the importance of selecting a proper tester community to ensure a representative evaluation of software.
@@ 2.5 Diversity in machine learning documentation frameworks
Recent documentation and reporting frameworks in the Machine Learning (ML) field, such as Datasheets for Datasets [13], Data Statements [54], and Data Cards [14], have become crucial to ensuring transparency, accountability, and fairness of these systems. These frameworks aim to capture the various aspects of data collection, model development, and evaluation while explicitly addressing biases, ethical considerations, and the societal impact of ML technologies. By incorporating diversity-related metadata into documentation, stakeholders can better assess whether ML systems fairly represent different demographic groups, mitigate potential biases, and adhere to inclusive algorithmic decision-making principles. Public regulatory initiatives such as the EU AI Act and the US AI Bill of Rights have started to adopt them, showing the increasing interest of public administrations in diversity.
A relevant contribution of these frameworks is their broader definition of the participants, which aligns with the perspective outlined in the previous subsections. They emphasize the role of individuals responsible for gathering and annotating training data, as well as those governing the project, highlighting the influence these groups can have on the final behavior of ML models. However, in the broader context of software projects, their focus is limited, as they concentrate almost exclusively on datasets while overlooking other critical assets such as code repositories and project management plans. Moreover, the range of relevant participants differs: whereas annotators are central in ML projects, developers, testers, and end users often have a greater impact in software development. Thus, although these frameworks underscore the importance of diversity and have attracted attention from public institutions, they fall short of capturing the complexity of teams in software projects.
@@ 3.1 Study method
The method we followed to analyze current practices on reporting diversity in OSS included
@@
The search for diversity documentation evidence has focused on five main groups extracted from the review of the previous section, and further presented in Section 4. Specifically: (A1) the development team, and whether they mention profile aspects about the developers; (A2) non-coding contributors, and whether they explain non-coding roles; (A3) tests with potential users, and whether they mention labor force or report platforms; (A4) usage context, including the target population and possible specific adaptations; and (A5) governance, including information on funders. Areas A1 and A2 are motivated by the works on diversity for developers and non-coding roles (cf. Sections 2.2 and 2.3), while A3 and A4 are inspired by the need for specifying human aspects (cf. Section 2.4). The works on ML documentation (cf. Section 2.5) prompted A5. For each area, we reported whether the OSS project included (or not) such information.
@@ 1. Does the text mention the development team explicitly?
Prompt and tool refinement. To perform the analysis and search for diversity documentation, we built a tool based on Large Language Models (LLMs) [^8] Tool’s repository: https://github.com/SOM-Research/Analyzer-Diversity-Card. In particular, we relied on GPT-4o-mini. The use of LLMs for analyzing text has demonstrated their effectiveness, for instance, for sentiment analysis [56, 57], text classification [58, 59], studying social interactions [60], and detecting ethical deviations [61]. In our context, the use of LLMs eliminates the need for extensive preprocessing, enabling a more streamlined and direct analysis of the text.
As a preliminary step before the automated analysis, we designed specific prompts for each of the five diversity areas considered (A1–A5). The prompts were initially drafted by one of the authors and iteratively refined in collaboration with two others, using a sample of 20 representative projects. This sample was used exclusively for prompt development and was not included in the final dataset. Each prompt follows a few-shot strategy and includes five complete examples, composed of an input text fragment and its corresponding output in structured JSON format. The examples cover a range of cases (i.e., from affirmative to negative and ambiguous) and illustrate different levels of detail regarding diversity-related aspects (e.g., team size, roles, gender, or geographic diversity). This structure is intended to guide the model both in terms of the content to be identified and the expected format of the response [^9] The full set of prompts is available at: https://github.com/SOM-Research/Analyzer-Diversity-Card/blob/main/src/classifier/prompts.py. Throughout the refinement process, both the wording and structure of the prompts were incresingly adjusted to promote consistency and reduce ambiguity in the model’s responses. Specific measures were also introduced to mitigate potential hallucinations, including the rephrasing of unclear expressions and the enforcement of strict output formatting in JSON, ensuring reproducible and interpretable results. A first evaluation on the 20 refinement projects continued until the model reached a performance level deemed satisfactory (above 90%). Listing 1 shows an excerpt of one of the five prompts used, in particular, the one addressing A1.
Validation. Once the prompts and methodology were stabilized, a formal validation was conducted on a random sample of 50 projects from the full dataset. For each project, the evidence extracted by the model across the five areas was manually reviewed and compared against the interpretation of the authors. The validation was carried out by two authors, who independently reviewed the outputs and discussed discrepancies until reaching full agreement. This process yielded a 92.54% agreement between the model’s output and human assessment, supporting the reliability of the tool for large-scale application.
Repository analysis. After the refinement and validation of the tool, we applied it to the full dataset of 1,000 repositories. For each project, the extracted non-code files were provided as input, and the prompts guided the LLM-based tool to identify the presence (or absence) of diversity information across the five defined areas (A1–A5). The classification results, together with the raw documentation data extracted from the projects, are openly available. [^10] Zenodo DOI: https://doi.org/10.5281/zenodo.14981689